OpenStudy (anonymous):

explain how you would find the exact value of 3 over the square root of 8 @primeralph

OpenStudy (zzr0ck3r):

3/sqrt(8) multiply top and bottom by srt(8) 3*sqrt(8)/8 = 3*sqrt(4)*sqrt(2)/8 = 3*4*sqrt(2) / 8 = 12*sqrt(2)/8 = 3*sqrt(2)/2

OpenStudy (primeralph):